OpenStudy (anonymous):

solve each equation and check 17x-15=12

OpenStudy (anonymous):

ok can you get the x by itself?

OpenStudy (anonymous):

first add the 15 to the 12

OpenStudy (alexandervonhumboldt2):

17x=12+15 x=(12+15)/17

OpenStudy (alexandervonhumboldt2):

x=27/17

OpenStudy (alexandervonhumboldt2):

17*27/17-15=12 27-15=12 12=12

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

you can reduce the 27/12 to 9/4

OpenStudy (anonymous):

your welcome :)